i think this is where is should be
for the polyhedron, use eulers formula to find the missing #s
faces?
edges 13
vertices 9
i know V-E+F=
9-13+f=

Answer by rothauserc(4718)

Euler's formula for geometry tells us that
F + V - E = 2
we are given edges = 13 and vertices = 9, therefore
F + 9 - 13 = 2
F + 9 = 15
F = 6
There are 6 faces